OpenStudy (anonymous):

Which is an equation in point-slope form for the given point and slope? Point: (5, 9); Slope: 2

OpenStudy (anonymous):

A. y – 5 = 2(x + 9) B. y – 9 = 2(x – 5) C. y + 9 = 2(x – 5) D. y – 9 = 2(x + 5)

OpenStudy (anonymous):

The answer is D

OpenStudy (anonymous):

Ty